OpenStudy (lgbasallote):

Note: This is NOT a question. This is a tutorial. What are the fundamental properties of logarithms? See comment below to see what! For beginners, see my tutorial. For more advanced people, scroll down through the comments to see the discussion of KingGeorge, satellite73, and amistre64 on other thingies that can be done with logarithms

OpenStudy (lgbasallote):

First of all, we need to understand the definition of logarithms. Definition of logarithms: \(\large \log_{a} b = x\) this means that x is the exponent that a must be raised to obtain b. Exponentially, we can rewrite it as \(\large a^{x} = b\) Now, on to the fundamental properties... Property 1: \(\large \log_{a} a = 1\) why? Simply because 1 is the only exponent you can raise a that it'll come out as itself. Example: \(\log_{3} 3 = 1\) Property 2: \(\large \log_{a} 1 = 0\) why? Because anything raised to zero is one. Of course, this property is only true if a is not equal to 1 nor zero. Why? Because if a is 1, there would be an infinite number of possible answers as 1 raised to anything will equal 1. On the other hand, \(0^0\) would not wield 1 either. \(0^0\) is an indeterminate form because it is a "gray area" between the properties "anything raised to zero equals one" and "zero raised to anything is zero". It satisfies both conditions, but it cannot be determined which is the answer for \(0^0\) Example: \(\log_{2} 1 = 0\) Property 3: \(\log_{a} a^x = x\) or \(a^{log_{a} x} = x\) These two are related. If we read the former, it is "what can we raise a to obtain \(a^x\)". Which obviously, the answer is x. For the latter, if we transform it to exponential form (which i am assuming you know how), you'll get \(\log_{a} x = \log_{a} x\) which shows that it is merely an identity property, the same with the former. Example: \(\log_{2} 2^3 = 3\) \(3^{log_{3} 2} = 2\) Product Law: \(\log_{a} x + \log_{a} y = \log_{a} xy\) note that the two addends MUST have THE SAME base. It merely states that if you add two logarithms with the same base, it is also equal to getting the product of your two powers (which in this case is x and y), then taking its logarithm to that same base (which in this case is a) Example: \(\log_{2} 3 + \log_{2} 5 = \log_{2} 15\) <---try inputting that in a scientific calculator and see if you get the same answers. *puke rainbow* Quotient Law: \(\large \log_{b}c - \log_{b}e = \log_{a} \frac{x}{y}\). This is somehow similar to the Product Law. In this property, it is stated that if it's the difference of two logaritms with the same base, it is also equal to taking the quotient of your two powers (which in this case is c and e) then taking its logarithm to that same base (which in this case is b). Example: \(\log_{3} 12 - \log_{3} 6 = \log_{3} 2\)<----again, you can try inputting it in the calculator to verify, but I'm sure you've ran out of rainbows to puke :D Power Law: \(\log_{a} b^x = x\log_{a} b\) Note that your base and power must NOT be the same. Otherwise, it is Property 3. This property merely states that if you have a power (which in this case is \(b^x\)) that also has an exponent, it is equal to solving for the logarithm of the base of the power (the power is \(b^x\) so the base of the power is b) to the base of the original logarithm (in this case, the base of our original logarithm is a) then multiplying it to the exponent of the power(which in this case is x). The definition I stated may sound confusing so a demonstration is in order. Example: \(\log_{2} 3^4 = 4 \log_{2} 3\) <---you know what to do..bring out those calculators and verify! *pukes pot of gold* Fun Fact: We can use the Power Rule to prove the first part of Propert 3. \(\log_{a} a^x\). If we apply Power Rule, it will become \(x\log_{a} a\). And what is \(\log_{a} a\)? It's 1! So, x(1) = x, thus proving that \(\log_{a} a^x = x\). Change-of-Base Formula. \(\Large \log_{a} b = \frac{\log_{c} b}{\log_{c} a}\). This, in my opinion, is the most confusing and the hardest to explain of all the fundamental properties. This property states that if you have a logarithm of b to the base a, it is also the equal to taking the logarithm of b to a certain base, then dividing it by the logarithm of a taken to the same base. You may ask, why is this essential? Why take two logarithms when you can only take 1? This property has many uses. Two of which is to transform bases. For example, you have \(\log_{3} 2\) but you need it as a base of two. You can use this property to change its base to 2. \(\Large \log_{3} 2 = \frac{\log{2} 2}{\log_{2} 3} = \frac{1}{\log_{2} 3}\) Now, it is in the base of 2. Another use? For usage of calculators. You cannot just merely input \(\log_{3} 2\) in a calculator, can you? (Actually some calculators can, but for the sake of explaining I'll asume you don't have that calculator). We can use this formula to do so! \(\log_{3} 2 = \frac{\log 2}{\log 3}\). You will learn later on what those invisible bases mean. For now, all you need to know is that \(\log 2\) and \(\log 3\) have the same "invisible bases", and that those logarithms with "invisible bases" are solveable by the common scientific calculator, and by using this property, you can solve for logarithms faster. Example: \(\log_{4} 2\) We are going to solve this in two ways. The first is the Change-of-Base Formula, the second is the application of the other fundamental properties to verify the answer. By C-O-B: \(\Large \log_{4} 2 = \frac{\log_{2} 2}{\log_{2} 4} = \frac{1}{\log_{2} 2^2} = \frac{1}{2}\) By Fundamental Properties: \(\Large \log_{4} 2 = \log_{4} 4^{\frac{1}{2}} = \frac{1}{2}\) \(\checkmark\)

OpenStudy (lgbasallote):

@KingGeorge

OpenStudy (anonymous):

Gj @lgbasallote ,Thats awesome :)

OpenStudy (lgbasallote):

thanks @Eyad :D

OpenStudy (kinggeorge):

Looking good.

OpenStudy (lgbasallote):

i had troubles with the change of base :/

OpenStudy (zepp):

Haha lol, I just learned logarithms a hour before this post, good job :D

OpenStudy (lgbasallote):

well now you know more :P btw Special mention to @zepp for showing me that tutorials can be funny by adding cliched puns :D

OpenStudy (zepp):

http://puu.sh/sXBu log22 ;)

OpenStudy (lgbasallote):

ARGHHHHHHHHH REWRITE!!! \LARGE \log_{3} 2 = \frac{\log_{2} 2}{\log_{2} 3}\)

OpenStudy (lgbasallote):

\(\LARGE \log_{3} 2 = \frac{\log_{2} 2}{\log_{2} 3}\)

OpenStudy (amistre64):

spose we simply define a set of functions that follow this simple rule: L(xy) = L(x) + L(y) and that any function that satisfies this condition is "Logarithmic"

OpenStudy (lgbasallote):

suppose? isn't it?

OpenStudy (amistre64):

based on this property; we should be able to define all the other properties that are so "mysteriously" logarithms i believe

OpenStudy (amistre64):

L(x) = L(x) ; why? L(x*1) = L(x) + L(1) = L(x) + 0 therefore L(1) = 0

OpenStudy (amistre64):

L(x^2) = L(x*x) = L(x) + L(x) = 2 L(x)

OpenStudy (lgbasallote):

i see...that is cool *_* they never teach you that :/

OpenStudy (amistre64):

Logarithms can be considered of without ever knowing what a Logarithm is :) just based on defining functions that have the property L(xy) = L(x) + L(y) L(x/x) = L(1) L(x* 1/x) = 0 L(x) + L(1/x) = 0 therefore: L(1/x) = -L(x)

OpenStudy (amistre64):

L(x/y) = L(x* 1/y) = L(x) + L(1/y) from above we can rewrite as L(x/y) = L(x) - L(y)

OpenStudy (amistre64):

http://www.cosmolearning.com/video-lectures/logarithms-without-exponents-10784/ im starting to love this guy :) his "hi" is a little creepy, but ehh....

OpenStudy (lgbasallote):

but i dont think it's easy to understand without the basic knowledge...i mean i was amused because it's a fascinating shortcut but i already know the fundamental properties...is it easy to learn withoout that knowledge?

OpenStudy (amistre64):

its simple enough to memorize; but if we really want to make any sense of it, we hav to delve deeper into it. The properties seem almost magical at first and fly in the face of what we are used to knowing as math. But, by knowing the fundamental basis of a Logarithmic function, and how its properties are defined and made rigorous thru calculus; we gain a deeper appreciation for them and the mystery fades

OpenStudy (lgbasallote):

that was deep...but based on what i comprehended i agree :)`

OpenStudy (amistre64):

about 13:15 in that video i linked he starts into what I have shared :)

OpenStudy (lgbasallote):

i wanna see the creepy hi >:))

OpenStudy (amistre64):

lol ive watched alot of his videos and they all start out with that creepy "hi" :)

OpenStudy (lgbasallote):

hahaha i keep rewatching it for that hi :p

OpenStudy (amistre64):

ive been thinking about how to go about defining a base to this system \[\frac{L(b)}{L(b)}=1;\ b\ne1\] as such we can define a base so be the value of 1/L(b) how do we come up with a logically consistent change of base formula from this? \[L_b(x)=\frac{L(x)}{L(b)},\ for\ reference \] spose we want to change the base from L(b) to L(b') . Wouldnt we just multiply by L(b)/L(b')? \[\frac{L(x)}{L(b)}*\frac{L(b)}{L(b')}=\frac{L(x)/L(b)}{L(b')/L(b)}=\frac{L(x)}{L(b')}\] \[L_b(x)= \frac{L_{b'}(x)}{L_{b'}(b)}=\frac{L(x)/L(b')}{L(b)/L(b')}=\frac{L(x)}{L(b)}\to\ L_b(x) \]

OpenStudy (lgbasallote):

hmm makes sense...

hero (hero):

@lgbasallote, while your at it, write a book :P

OpenStudy (lgbasallote):

am working on it :P hahaha

OpenStudy (anonymous):

you can show all these properties belong to \[L(x)=\int_1^x\frac{1}{t}dt\] directly

OpenStudy (lgbasallote):

really? the properties are in integral? :O

OpenStudy (anonymous):

for example, you can show that \[L(ab)=L(a)+L(b)\] by showing that \[\int_1^{ax}\frac{1}{t}dt=\int_1^x\frac{1}{t}dt +\int_1^a\frac{1}{t}dt\]

OpenStudy (anonymous):

yes they are definitely in the integral because the most precise definition of \(\log(x)\) is \[\log(x)=\int_1^x\frac{1}{t}dt\]

OpenStudy (lgbasallote):

ommigosh o.O

OpenStudy (kinggeorge):

I'd also like to chime in saying that logarithms are not restricted to the real numbers. In fact, they're commonly used in cyclic groups of prime order as well. Although you only really see this application in number theory, and/or cryptography.

OpenStudy (anonymous):

as in \(\mathbb{Z_p}\) we have the "discrete logarithm"

OpenStudy (kinggeorge):

Without which the Diffie-Hellman Key Exchange Protocol would be much more difficult, Along with El-Gamal Public Key Cryptosystem.

OpenStudy (anonymous):

and dont forget \(\log(z)=\ln(|z|)+i\theta\)

OpenStudy (anonymous):

And one more property \[\huge a^{\log_b c }= c^{\log_b a}\]

OpenStudy (unklerhaukus):

Tell me more about logarithms of complex numbers \[\log(z)=\ln(|z|)+iθ\] Where does this formula come from ? How come the base changes?

OpenStudy (anonymous):

@UnkleRhaukus z can be expressed as\[ z = |z|e^{i(\theta)}\] Now apply log on both sides and get your required answer

OpenStudy (anonymous):

How come the base changes? --->Actually it should be \[\ln(z)=\ln(|z|)+i\theta\]

OpenStudy (unklerhaukus):

\[~~~~~~z=|z|e^{iθ}\]\[\qquad\downarrow\]\[\{\text{taking the natural logarithm}\}\]\[\qquad\downarrow\]\[\ln(z)=\ln|z|e^{iθ}\]\[=\ln |z|+i\theta\ln e\]\[=\ln|z|+i\theta\]

OpenStudy (anonymous):

Exactly @UnkleRhaukus

OpenStudy (unklerhaukus):

Is the wikipedia page also incorrect @shivam_bhalla ? http://en.wikipedia.org/wiki/Complex_logarithm

OpenStudy (kinggeorge):

From my reading of the wiki page, the complex log doesn't have a specified base. The log is defined in the manner that makes the most sense. It's just the component \(\ln(r)\) that has an actual base.

OpenStudy (kinggeorge):

Additionally, this can't be seen as an ordinary logarithm since, the property \[\log(ab)=\log(a)+\log(b)\]fails.

OpenStudy (anonymous):

Yes We are specifically considering to the base e while taking log. @KingGeorge is right :)

Directrix (directrix):

Practice Question. What real number values of x satisfy the following logarithmic equation? See attachment.

OpenStudy (zepp):

\(\huge (x+1)^{\log_{2}(x+1)^8}=3\) \(\huge \log_{x+1}3=\log_2 (x+1)^8\) Let's say x+1 = a Then \(\huge \log_{a}3=8\log_2 a\) \(\huge 8 = \frac{log_a 3}{log_2 a}\) And then I'm stuck :(

OpenStudy (zepp):

@KingGeorge

OpenStudy (anonymous):

It's easy \[8(\log_2(x+1))^2 = 8\] \[(\log_2(x+1))^2=1\]

OpenStudy (anonymous):

I am sorry but I just shortened my steps to a great extent :D

OpenStudy (zepp):

lol where does that square come from? o.o

OpenStudy (kinggeorge):

\[\Large 8 = \frac{log_a 3}{log_2 a}=\frac{\frac{log_2 3}{log_2a}}{log_2 a}\] \[\Large=\frac{log_2 3}{\left(log_2a\right)^2}\]

OpenStudy (kinggeorge):

Simplify a bit more and you get what @shivam_bhalla got.

OpenStudy (anonymous):

OK :D. Looks like I should give you a hint Use this property \[\log_a (b^c) = c \log_a (b)\]

OpenStudy (zepp):

So \(\huge 8=\frac{log_2 3}{(log_2 a)^2}\) \(\huge 8(log_2 a)^2 = log_2 3\) ?

OpenStudy (zepp):

8(log2(x+1))2=8 I can't get the 8 you have on the right side D;

OpenStudy (anonymous):

LOL. Let me show it to you in another way :D \[\large \log_2{(x+1)^{\log_2{(x+1)}^8}}=8\] Now using the property I mentioned before, we get \[\large(\log_2{(x+1)}^8) (\log_2{(x+1)})=8\] Now again applying the property, \[\large8(\log_2{(x+1)}) (\log_2{(x+1)})=8\] which is \[\large8(\log_2{(x+1)}^2)=8\]

OpenStudy (anonymous):

Sorry Typo in the last step. It should be \[\large8(\log_2{(x+1)})^2=8\]

OpenStudy (zepp):

OH! I see I see :D Thanks @shivam_bhalla! And @KingGeorge Isn't \(\log_2{8} =3\) ? \(\log_2(3)\) gives me 1.58496..

OpenStudy (kinggeorge):

I realized that -.- I am disappoint.

OpenStudy (anonymous):

@zepp , wasn't it easy. :D Now take two cases Case 1-> \[\large \log_2{(x+1)}=1\] Case 2-> \[\large \log_2{(x+1)}=-1\] And solve. But One doubt I have is whether you have take either union or intersection of the answers. @KingGeorge , @zepp , Any help in this part :)

OpenStudy (anonymous):

@KingGeorge , Check again :D You are going wrong.

OpenStudy (kinggeorge):

Ignore my ramblings.

OpenStudy (zepp):

x = 1 or -0.5? Now I'm trying to understand why my method didn't work :|

OpenStudy (anonymous):

Ok. The mistake you are making is that it should be \[\large \log_a( \frac{b}{c}) = \log_a b - \log_a c\]

OpenStudy (anonymous):

But in your case, you have taken it as \[\large \frac{\log_a( b)}{\log_a (c)} \neq \log_a b - \log_a c\]

OpenStudy (kinggeorge):

I'm good at this. I promise! Except when I do it around other people :(

OpenStudy (anonymous):

LOL @KingGeorge . It happens many a time with me too. But for me, this happens with topic :integration :D

OpenStudy (zepp):

meh, I still don't see my mistake :|

OpenStudy (anonymous):

OK. :D Let me have a check :)

OpenStudy (zepp):

Ah thanks :D

OpenStudy (anonymous):

Your first step itself is a mistake. It should be \[\huge (x+1)^{\log_{2}(x+1)^8}=2^8\]

OpenStudy (zepp):

Oh, right :x

OpenStudy (zepp):

Herp derp, I probably thought that \(log_2 y = 8\) y = 3

OpenStudy (anonymous):

No problem . Remember \[\large \log_a b = c\] Then \[\large b = a^c\]

OpenStudy (zepp):

Haha, I know those properties, it's probably because of the 2 and the 8 :P Anyway, thanks guys, it's quite surprising how much we can learn from 1 question! ;D

OpenStudy (anonymous):

Yes. Thanks to @Directrix for the question and both of you @KingGeorge @zepp :)

OpenStudy (zepp):

Yeah, thanks to @Directrix and @KingGeorge :)

OpenStudy (zepp):

and @shivam_bhalla, of course!

Directrix (directrix):

Another Log Practice Problem For This Tutorial: See Attachment.

OpenStudy (zepp):

Kellogg! :D lol let me do it :)

OpenStudy (zepp):

I guess the first step would be \(\large \log_k 16 = \log_2 5\) Then \(\huge \frac{\log_2 16}{log_2 k}=log_2 5\) \(\huge \frac{4}{log_2 5} = log_2 k\) \(\large k \approx 3.300549\) Plug it in the question and the answer would be 625.